(t/3 -2)=(3t/8)
how do you solve?

Answer by checkley77(12844)

(t/3 -2)=(3t/8)
t/3-2=3t/8 subtract 3t/8
t/3-3t/8=2 combine the fractions by finding the common denominator (3*8=24)
(8*t-3*3t)/24=2 multiply the numerator terms.
(8t-9t)/24=2
-t/24=2 cross multiply.
-t=48
t=-48 ans.
Proof:
-48/3-2=3*-48/8
-16-2=-144/8
-18=-18